OpenStudy (anonymous):

Any help is nice. Eliminate the perimeter x=1/2 t, y=(3t^3) - 1

OpenStudy (anonymous):

\[x=\frac{ 1 }{ 2 }t , y=3t^3 - 1\]

OpenStudy (mathmale):

Ross: Please look up the word "eliminate." Are you sure this is the word you want to use here?

OpenStudy (anonymous):

Honestly its not the word choice I would go with but the problem given uses that exact phrasing

OpenStudy (mathmale):

Look up "perimeter" also. sure you want to use that word?

OpenStudy (triciaal):

i think the wrong word is perimeter ;perhaps should be parameter

OpenStudy (mathmale):

Ross, would you want to eliminate perimeters for a living? Think about it.

OpenStudy (anonymous):

Parameter. Oops. I read that wrong Its eliminate the parameter. didnt notice till triciaal said something

OpenStudy (mathmale):

I seriously hope you will actually look up the words in question. Yes, it's parameter, not perimeter. Makes a huge difference. So, you are to eliminate the parameter here. What is the parameter? Identify it.

OpenStudy (mathmale):

Hint: you are given a formula for x and another one for y. Both x and y are functions of what parameter?

OpenStudy (anonymous):

Sorry was tabbed out for second. If im correct the goal is to combine the two equations while solving for t

OpenStudy (mathmale):

The goal is to eliminate the parameter, t, so that y ou end up with a single equation giving y as a function of x or x as a function of y.

OpenStudy (mathmale):

Of the 2 formulas given you, one for x(t) and one for y(t), which would be the easier to solve for t?

OpenStudy (anonymous):

x for t. x= 1/2t that would become 2x = t?

OpenStudy (mathmale):

You began with x=(1/2)t. Please substitute your 2x=t into that first equation as a check.

OpenStudy (anonymous):

x=(1/2)t x=(1/2)2x x=x

OpenStudy (mathmale):

OK. Let's go with that. Always good to check. We want to eliminate the parameter, t. What's our next step, using your t=2x?

OpenStudy (anonymous):

for each value of t in the equation y=(3t^3) - 1 substitute in 2x? y = (3 * 2x ^3) - 1 y = (6x^3) - 1

OpenStudy (mathmale):

You are given the equation y=(3t^3)-1. simply throw out the "t" and write (2x) in its place. We have to be really careful where we place those parentheses.

OpenStudy (mathmale):

Think of that y=(3t^3)-1 as\[y=(3)(t)^3-1.\]

OpenStudy (mathmale):

I'm suggesting that you throw out the t and replace it with 2x.

OpenStudy (mathmale):

Ross?

OpenStudy (anonymous):

Sorry was reading that over. Isnt that what I did? Maybe i did the order wrong? \[y=(3)(2x)^3 -1\] so then 8x^3 instead?

OpenStudy (mathmale):

Yes, now you have it right. y = 3 (8x^3) - 1. Can you simplify that further?

OpenStudy (anonymous):

\[y=24x^3 - 1\] Then that would be the end result?

OpenStudy (mathmale):

Yes. So, Ross , you now have an equation for y in terms of x alone, and have succeeded in eliminating the parameter, t. Very good!

OpenStudy (anonymous):

Awesome. Thank you Mathmale. Ill close the question now. That helped alot

OpenStudy (mathmale):

so glad to hear it! Thx for your perseverance. Merry Christmas!